OpenStudy (peterg2001):

help will medal Subtract 2 16/21 - (-8 5/21) Reduce, if possible.

OpenStudy (anonymous):

2 16/21 = 58/21 -8 5/21 = -173/21 58/21 - (-173/21) 58/21 + 173/21 231/21 11 C

OpenStudy (anonymous):

I think It would be 11 cause (2 16/21) - (-(8 5/21)) = 11

OpenStudy (anonymous):

@thatonechick1 is right